Assume that a company operates a fleet of limousines. If a limo is driven 80,000 miles during a year its average operating cost is 25 cents per mile. If a limo is driven only 60,000 miles during the year, its average operating cost is 30 cents per mile. Using the high-low method, what is the estimated fixed cost per year?
Answer #1

Total cost at 80,000 miles=(80,000*0.25)=$20,000

Total cost at 60,000 miles=(60,000*0.3)=$18000

Variable cost per mile=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(20,000-18000)/(80,000-60,000)

=$0.1 per mile

Total fixed cost=20,000-(80,000*0.1)

=$12000